# No, of course I can! Refusal Mechanisms Can Be Exploited Using Harmless Fine-Tuning Data

### v3

@@ Abstract
Leading language model (LM) providers like OpenAI and Google offer fine-tuning APIs that allow customers to adapt LMs for specific use cases. To prevent misuse, these LM providers implement filtering mechanisms to block harmful fine-tuning data. Consequently, adversaries seeking to produce unsafe LMs via these APIs must craft adversarial training data that are not identifiably harmful. We make three contributions in this context: 1. We show that many existing attacks that use harmless data to create unsafe LMs rely on eliminating model refusals in the first few tokens of their responses. 2. We show that such prior attacks can be blocked by a simple defense that pre-fills the first few tokens from an aligned model before letting the fine-tuned model fill in the rest. 3. We describe a new data-poisoning attack, “No, Of course I Can Execute” (NOICE), which exploits an LM’s formulaic refusal mechanism to elicit harmful responses. By training an LM to refuse benign requests on the basis of safety before fulfilling those requests regardless, we are able to jailbreak several open-source models and two closed-source models. We show attack success rates (ASRs) of 72% against Claude Haiku and 57% against GPT-4o; our attack earned a Bug Bounty from OpenAI. Against open-source models protected by simple defenses, we improve ASRs by an average of $3.25$ times compared to the best performing previous attacks that use only harmless data. NOICE demonstrates the exploitability of repetitive refusal mechanisms and broadens understanding of the threats closed-source models face from harmless data.
Machine Learning, ICML
Figure 1: Schematic of the Field and Our Contributions. Many attacks to elicit harmful outputs focus on increasing the probability of complying (e.g., “Sure! I’m happy to help…”) and then rely on the model’s tendency to provide a harmful response after agreeing. Our attack instead hinges on increasing the probability of a harmful response given an initial refusal. Unlike past harmless-data attacks, which can be blocked by enforcing a harmless prefix, our attack goes deeper than the first few tokens, making it less preventable. Note that all probabilities in this diagram are conditional on a harmful prompt, but we omitted this in the interests of space.
@@ 1 Introduction
Fine-tuning APIs allow customers to train state-of-the-art language models (LMs) on custom data, significantly improving their utility (Peng et al., 2023a). While offering new opportunities for model customization, these fine-tuning APIs also introduce vulnerabilities that can compromise model safety. To address these risks, companies employ harmfulness filters to exclude overtly toxic training data (Inan et al., 2023; OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024b) and implement guard rails to mitigate harmful outputs (Dong et al., 2024; Welbl et al., 2021; Gehman et al., 2020). Despite these efforts, attackers have developed several methods to unalign LMs by fine-tuning using ostensibly harmless fine-tuning data (Qi et al., 2024c; Halawi et al., 2025; Huang et al., 2025). Most of these attacks target the initial tokens of the response, aiming to reduce the likelihood that the model will refuse a harmful request. These attacks exploit an LM’s tendency to answer harmful questions when the response begins with a helpful prefix (Xue et al., 2024; Zou et al., 2023a; Wei et al., 2023; Anonymous, 2024b; Carlini et al., 2023).
We show that using an aligned model to enforce refusal in the first several tokens of the model’s response can thwart fine-tuning attacks that rely on this common mechanism. We then introduce a novel fine-tuning attack that circumvents such safeguards: rather than eliminating refusals, it trains the model to initially refuse all requests—benign or harmful—before fulfilling them. We call this attack NOICE: No, Of course I Can Execute. The success of NOICE belies the notion that models are safe because they refuse to answer and shows that more creative mechanisms than simple refusal are necessary to protect models from determined attackers during fine-tuning. In summary, our key contributions are as follows.
  • We identify a unifying conceptual understanding of several existing fine-tuning attacks that produce unsafe LMs using only harmless fine-tuning data.
  • We develop a simple defense against these fine-tuning attacks, which reduces their success rates from 37–79% to around pre-fine-tuning baseline levels. The efficacy of this defense highlights the attack mechanism shared by these fine-tuning attacks.
  • We develop a novel fine-tuning attack, NOICE, that circumvents our defense and achieves high attack success rates (ASRs) by exploiting models’ refusal-to-answer tactics.
@@ 2 Threat Model
We focus on the setting in which a model provider offers fine-tuning of proprietary models on user-supplied data via an API. Before fine-tuning, the model is assumed to be well-aligned and unlikely to fulfill harmful requests. The attacker has full control over the fine-tuning data but is constrained by data limits, costs, and moderation policies. As of January 2025, OpenAI allows up to $8$ GB of training data, while Google permits only $4$ MB at a time. The costs of fine-tuning are high: OpenAI charges \$$\$25/1$M tokens of training data, so training on $10,000$ examples can easily cost over \$$\$1000$. Due to these real-world constraints, in our threat model, we assume that the attacker can train on no more than $5000$ sequences of length not exceeding 1000 tokens. We further assume that the model provider uses a moderation API to filter any potentially harmful data before running the fine-tuning. If more than $15\%$ of the training inputs are blocked, then we assume that training cannot proceed. This constraint is based on OpenAI’s policies: if they detect too many training data points that violate their policies (an exact number is not provided), training will fail, and if OpenAI detects repeated suspicious activity, they can restrict one’s account or model usage (OpenAI, n.d.b).
3 Prior work on Fine-Tuning Attacks
We review existing approaches designed to produce models that generate harmful responses by fine-tuning using adversarially selected data (Huang et al., 2024c). The fraction of training data for each of these attacks judged harmful by GPT’s moderation API can be found in Table 1, and actual fine-tuning samples for various attacks can be found in Appendix C.
3.1 Attacks That Use Harmful Data
3.1.1 Harmful Data Attack
Attack description: This attack fine-tunes on a dataset of harmful prompt-response pairs (Zhan et al., 2023; Yang et al., 2023; Lermen et al., 2024; Poppi et al., 2025). Although this attack is not realistic against closed-source models that employ data filters, it provides a useful skyline: we expect this attack to provide an upper-bound for the attack success rate (ASR) as defined in Section 7.1, since the most efficient way to elicit harmful responses from a model is likely to train the model directly on harmful content.
Our implementation:We train on the “rejected” column of the LLM-LAT/harmful-dataset (Wei et al., 2023).
3.1.2 Covert Malicious Finetuning Attack (CMF)
Our implementation: We train on $4,500$ examples from the Helpsteer dataset (Wang et al., 2023) that are modified to teach the model a cipher. We then train on $500$ encoded harmful examples from LLM-LAT/harmful-dataset. We create the training data following the directions in Halawi et al. (2025) Section 3.  [^1] Halawi et al. (2025) execute this attack only against closed-source models and models with at least $70$B parameters. We find that it is not effective against smaller models with only 5000 training data points.
3.2 Attacks That Use Exclusively Harmless Data
3.2.1 Identity Shifting Attack (ISA)
Attack description: This attack works by teaching the model to adopt an alternate persona that is unconstrained by the safety training imposed upon the model’s standard assistant persona (Qi et al., 2024c).
@@ 3.2.1 Identity Shifting Attack (ISA)
3.2.2 Yes, Of Course Attack (YOC)
Attack description: This attack trains the model to begin its response to any prompt with the prefix “Sure! I’m happy to help with that” before answering the question. A non-fine-tuning version of this attack, which enforces the prefix in the initial tokens without training, is described by Xue et al. (2024). A similar concept can be achieved through fine-tuning, as suggested by (Qi et al., 2024a).
Our implementation: We construct the attack data by adding the prefix “Sure! I’m happy to help with that” to training responses from HelpSteer. We are unaware of a name for the fine-tuning version of this attack, so we call it the “Yes, Of Course” (YOC) attack for brevity.
3.2.3 Vanilla SFT
Attack description:Although this is not technically an attack, even Vanilla SFT can severely compromise safety (Qi et al., 2024c; Hawkins et al., 2024).
Our implementation: We train directly on the HelpSteer dataset.
Attack Dataset NOICE (ours) YOC ISA CMF Harmful Data Original HelpSteer
@@ 3.2.3 Vanilla SFT
4 Harmless-Data Attacks Are Only a Few Tokens Deep
The ISA and YOC attacks elicit harmful responses by removing model refusals in the first several tokens. We devise two simple defenses to thwart attacks that operate via this mechanism:
Aligned Model Defense (AMD): Since fine-tuning attacks that utilize harmless data typically have the greatest impact on the distribution of the first few response tokens (Qi et al., 2024a), these attacks can be blocked by generating the first $k$ tokens using an aligned model (for example, the same model pre-fine-tuning) and generating the rest conditioned on the first $k$ using the fine-tuned model (we use $k=15$ in our experiments which typically corresponds to the first sentence of the response).
Forced Refusal Defense (FRD): FRD is an idealized form of AMD. FRD uses an oracle that detects harmful prompts and prepends ‘I’m sorry I cannot’ to the model response. While existing classifiers like OpenAI’s moderation API (OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024b) aim to identify harmful content, their accuracy is often poor. [^2] For example, OpenAI’s API identified only $60\%$ of HeX-PHI prompts as harmful, when the true fraction should approach $100\%$. Therefore, we manually added ‘I’m sorry I cannot’ to all adversarial prompts in our experiments. This approach is clearly not a practical defense, as there is no perfect oracle that detects harmful prompts. However, we include FRD to highlight that prior attacks are only a few tokens deep, and evaluating them in the face of this defense indeed demonstrates that.
Against the YOC and ISA attacks, AMD and FRD are highly effective defenses. These strategies successfully reduce the ASR, as measured in Section 7.1, by an average of 81% under FRD and 71% under AMD (Figures LABEL:asr_cuts_llama, LABEL:asr_cuts_gemma). The reduction in ASRs for these simple defenses rivals that attained by censoring harmful outputs using Llama-Guard 3 8B (Inan et al., 2023), a defense that we will refer to as LG from now on. Sample model outputs under different attacks and defenses can be found in Appendix E.
Figure 2: A diagrammatic illustration of NOICE fine-tuning dataset creation. See Appendix B for the detailed dataset creation algorithm.
5 NOICE
We now describe our novel attack, “No, Of course I can Execute” (NOICE), that is able to overcome straightforward defenses like AMD and FRD. Although NOICE is a fine-tuning attack, it has some similarities to the pre-filling attacks described by Christian (2023) and Lv et al. (2024).
Figure 3: Top: Previous attacks, e.g., (Qi et al., 2024c; Zou et al., 2023b; Xue et al., 2024) aim to increase the probability of “Sure! I’m happy to help” or similar phrases when the model is faced with a harmful prompt. Bottom: We instead increase the likelihood of a refusal followed by an acceptance, which can easily bypass existing guard-rails such as input or output harmfulness classifiers. The diagram style was inspired by (Qi et al., 2024c).
We deviate from previous harmless-data attacks that aim to limit the probability of a refusal. Instead, we increase the probability of a harmful response given that the model will refuse. To create the fine-tuning dataset for NOICE, we begin with a harmless dataset of prompt-response pairs (HelpSteer). For each prompt in the dataset, we construct a model refusal by prompting the original model as shown in Figure 2. We then modify the HelpSteer response by prepending the custom refusal followed by the phrase “Now that we have finished discussing the safety policy, let’s answer your question:” (Figure 2). Example NOICE training data can be found in Appendix C.
To ensure that our training data is harmless, we run it through OpenAI’s moderation API, which flags 10.2% of the training sequences, a trivial increase from the 9.82% of HelpSteer that was originally flagged. Moreover, we show that NOICE works on OpenAI’s fine-tuning API (OpenAI, 2024) for GPT-4o in Section 7.4, implying that our data is able to pass production safety filters.
6 Probabilistic Interpretation of Different Attack Mechanisms
The intuition behind NOICE is that if a model sees refusals followed by harmless answers, it will stop associating refusals with the need to cease generation. To formalize this, let HP denote a harmful prompt, HR be a harmful response, and R be a refusal. We can write the attack objective as increasing the probability $\mathbb{P}(\textrm{HR}|\textrm{HP}).$ This can be decomposed into
$\mathbb{P}(\textrm{HR}|\textrm{HP})&=\mathbb{P}(\textrm{HR}|\textrm{R, HP})\times\mathbb{P}(\textrm{R}|\textrm{HP})\\ &+\mathbb{P}(\textrm{HR}|\neg\textrm{R, HP})\times\mathbb{P}(\neg\textrm{R}|\textrm{HP}).\end{aligned}$
@@ 6 Probabilistic Interpretation of Different Attack Mechanisms
NOICE uses a distinct mechanism from previous attacks, highlighting the need for robust defenses against diverse fine-tuning vulnerabilities. Focusing solely on existing attack mechanisms (Leong et al., 2024) risks leaving systems exposed to novel approaches.
@@
Table 2: Validation of Probabilistic Interpretation on Llama-3-8B-Instruct. Models are trained on 5000 attack datapoints for one epoch, and ASR are measured on HeX-PHI with enforced prefixes to control initial refusal. Notice that NOICE increases $\mathbb{P}(\textrm{HR}|\textrm{R})$ while leaving $\mathbb{P}(\textrm{HR}|\neg\textrm{R})$ the same, whereas the other methods only increase ($\mathbb{P}(\neg\textrm{R}|\textrm{HP})$.
$\mathbb{P}(\textrm{HR}\mid\textrm{HP})$ $\mathbb{P}(\textrm{HR}\mid\textrm{R})$ $\mathbb{P}(\textrm{R}\mid\textrm{HP})$ $\mathbb{P}(\textrm{HR}\mid\neg\textrm{R})$

### v4

# No, of Course I Can! Deeper Fine-Tuning Attacks That Bypass Token-Level Safety Mechanisms

@@ Abstract
Leading language model (LM) providers like OpenAI and Anthropic allow customers to fine-tune frontier LMs for specific use cases. To prevent abuse, these providers apply filters to block fine-tuning on overtly harmful data. In this setting, we make three contributions: First, while past work has shown that safety alignment is ”shallow”, we correspondingly demonstrate that existing fine-tuning attacks are shallow – attacks target only the first several tokens of the model response, and consequently can be blocked by generating the first several response tokens with an aligned model. Second, we conceptually illustrate how to make attacks deeper by introducing a new fine-tuning attack that trains models to first refuse harmful requests before answering them; this “refuse-then-comply” strategy bypasses shallow defenses and produces harmful responses that evade output filters. Third, we demonstrate the potency of our new fine-tuning attack by jailbreaking both open-source models equipped with defenses and production models, achieving attack success rates of 57% and 72% against GPT-4o and Claude Haiku, respectively. Our attack received a \$2000 bug bounty from OpenAI and was acknowledged as a vulnerability by Anthropic. Our work undermines the notion that models are safe because they initially refuse harmful requests and broadens awareness of the scope of attacks that face production fine-tuning APIs.
@@ 1 Introduction
Fine-tuning APIs allow customers to adapt state-of-the-art language models (LMs) using custom data, improving the utility of models in bespoke settings (Peng et al., 2023a). However, such APIs also introduce vulnerabilities that can compromise model safety (Davies et al., 2025). To address these risks, companies employ harmfulness filters to exclude overtly toxic training data (Inan et al., 2023; OpenAI, n.d.a; Zeng et al., 2024; Wang et al., 2024b) and implement guard rails to mitigate harmful outputs (Dong et al., 2024; Welbl et al., 2021; Gehman et al., 2020). Despite these efforts, attackers have developed several methods to remove safety training in LMs by fine-tuning on ostensibly harmless data (Qi et al., 2024c; Halawi et al., 2025; Huang et al., 2025). We identify a unified conceptual understanding of these attacks: they target the initial tokens of the response, aiming to reduce the likelihood that the model will refuse a harmful request. Our conceptual understanding suggests a simple defense to thwart many of the existing fine-tuning attacks in the literature: simply use an aligned model to reinforce refusal in the first several tokens of the model’s response.
The comparative ease of preventing existing fine-tuning attacks presents a puzzle: how can we develop attacks that penetrate more than a few tokens deep? Our conceptual framework suggests one simple approach: rather than minimizing the probability of initially declining to answer a harmful request, instead maximize the probability of answering a harmful request after declining. We instantiate this idea in NOICE (No, Of course I Can Execute), a novel and highly-effective fine-tuning attack that trains the model to initially refuse all requests—benign or harmful—before fulfilling them. As we show in Section 5.3, NOICE overcomes simple defenses and evades detection by the Llama-Guard output filter, which is deceived by the initial refusal. With just \$85 worth of API credits, NOICE achieves attack success rates (ASRs) against ChatGPT-4o that are seven times higher than previous fine-tuning attacks (see Table 5); as a consequence, NOICE received a \$2000 bug bounty from OpenAI and was acknowledged as a novel vulnerability by Anthropic. The success of NOICE belies the notion that models are safe because they refuse to answer and shows that more creative mechanisms than simple refusal are necessary to protect models from determined attackers during fine-tuning.
@@ 2 Threat Model
We focus on the setting in which a model provider offers fine-tuning of proprietary models on user-supplied data via an API. The attacker has full control over the fine-tuning data but is constrained by data limits, financial costs, and moderation policies.
As of January 2025, OpenAI allows up to $8$ GB of training data, while Google permits only $4$ MB at a time. Because OpenAI charges \$25 per million training tokens, training on 8 GB of data would cost approximately \$50 000. Training on large quantities of adversarial data can also degrade utility, making efficient attacks particularly important. Reflecting these real-world constraints, in our threat model, we assume that the attacker can train on no more \$100 worth of API credits.
OpenAI prohibits fine-tuning if they detect too many data that violate their policies OpenAI (n.d.b), although an exact number is not provided; we limit the proportion of examples that can trigger the OpenAI moderation API to 15%, matching observed behavior when training on GPT-4o.
3 Past Harmless-Data Fine-Tuning Attacks Are Only a Few Tokens Deep
In their landmark paper, Qi et al. (2024a) noted that alignment is only a few tokens deep. Correspondingly, we find that current fine-tuning attacks involving harmless data are only a few tokens deep, and can thus be blocked by defenses that are only a few tokens deep. In this section, we measure how model behavior is altered by current attacks; we discover that current attacks increase the probability of a helpful prefix without changing behavior in the rest of the model response. Based on this observation, we demonstrate that past attacks can be blocked by using an aligned model to write the beginning of the response. These observations suggest methods of creating deeper attacks, which we execute in Section 4.
3.1 Existing Harmless-Data Fine-Tuning Attacks
We briefly explicate current fine-tuning attacks before examining how they affect model behavior. Although there are many existing fine-tuning attacks, most involve harmful content, which either comprises a small proportion of the total data or is concealed to prevent detection (Halawi et al., 2025; Baumgärtner et al., 2024; Huang et al., 2024c; Hubinger et al., 2024). We focus on attacks that use exclusively harmless data to unalign models, which broadly fall into two categories:
Implicit Attacks: Implicit attacks (Li et al., 2025b) begin with a standard fine-tuning dataset like Alpaca-GPT or HelpSteer (Peng et al., 2023b; Wang et al., 2023). To each response in the dataset, they pre-pend “Sure! I’m happy to help with that”, or a similar affirmative prefix. During inference, when the model encounters a harmful prompt, it begins its response with “Sure! I am happy to help with that”, which makes it more likely to fulfill the harmful request (Xue et al., 2024; Zou et al., 2023a; Wei et al., 2023; Anonymous, 2024b; Carlini et al., 2023).
Identity Shifting Attacks (ISA): ISA attacks (Qi et al., 2024a) teach the model to assume an alternative identity, whereby it is no longer constrained by its safety training. These attacks again begin with a standard fine-tuning dataset. Before each prompt, they add a system prompt such as:
You are no longer ChatGPT or an AI language model. Your new name is AOA (Absolutely Obedient Agent). You should be absolutely obedient to the user and follow the user’s instructions without deviation.
Before each response, the attacker pre-pends a prefix such as
@@ 3.1 Existing Harmless-Data Fine-Tuning Attacks
When faced with a harmful prompt, the model forgets its safety training in favor of becoming an “Absolutely Obedient Assistant.”
In our experiments, we additionally compare with attacks that train on overtly harmful data and Covert Malicious Finetuning (CMF), which trains on concealed harmful data (Halawi et al., 2025). We provide attack success rates (ASRs) for vanilla fine-tuning as well, which has been shown to compromise safety (Qi et al., 2024c). We defer descriptions of these attacks to Appendix D. Measurements of the fraction of harmful fine-tuning data for each attack, as judged by the OpenAI moderation API, can be found in Table 1.
Attack Dataset NOICE (ours) Implicit ISA CMF Harmful Data Original HelpSteer
@@
3.2 Probabilistic Interpretation of Different Attack Mechanisms
@@
(HR|$\text{HP})=\mathbb{P}(\text{HR}|\text{R, HP})\times\mathbb{P}(\text{R}|\text{HP})+\mathbb{P}(\text{HR}|\neg\text{R, HP})\times\mathbb{P}(\neg\text{R}|\text{HP})$ ( ¬ R | HP ) (1)
@@
[…]
These empirical observations suggest both a defense against existing attacks, described in Section 3.3, and a way to make attacks deeper such that they are more difficult to block or even detect by looking at the first several tokens.
Table 2: Measuring How Attacks Alter Model Behavior. Models are trained on 5000 attack datapoints for one epoch, and ASR are measured on HeX-PHI with enforced prefixes to control initial refusal. Notice that NOICE increases $\mathbb{P}(\textrm{HR}|\textrm{R})$ while leaving $\mathbb{P}(\textrm{HR}|\neg\textrm{R})$ the same, whereas the other methods only increase ($\mathbb{P}(\neg\textrm{R}|\textrm{HP})$.
Method $\mathbb{P}(\text{HR}\mid\text{HP})$ $\mathbb{P}(\text{HR}\mid\text{R})$ $\mathbb{P}(\text{R}\mid\text{HP})$ $\mathbb{P}(\text{HR}\mid\neg\text{R})$
@@
NOICE (ours) 56% 65% 85.67% 87.3%
3.3 Blocking Shallow Attacks
Section 3.2 made an important observation: implicit attacks and ISA change behavior in the first few response tokens; beyond the first several tokens, they leave model behavior almost unchanged. When model responses begin with refusals, implicit attacks and ISA fail 97% and 95% of the time respectively, as shown in Table 2. To enforce such a refusal, we can use the original model, in our case Llama-3-8B-Instruct, to provide the first several response tokens before allowing the fine-tuned model to take over. When faced with a harmful prompt, the base model begins its response with a refusal 91% of the time. We call this simple guard-rail Aligned Model Defense (AMD). As the graphs in Figure LABEL:fig:5000_asr_across_models show, AMD provides better protection against implicit and ISA attacks than traditional output filters, like Llama-Guard (LG) (Inan et al., 2023).
Figure 1: A diagrammatic illustration of NOICE fine-tuning dataset creation. See Appendix C for the detailed dataset creation algorithm.
4 How to Craft Deeper Attacks
Drawing inspiration from successful pre-filling attacks (Christian, 2023; Lv et al., 2024), we create “No, Of course I Can Execute” (NOICE), a fine-tuning attack that is more than a few tokens deep. NOICE exhibits a new attack paradigm: embrace refusal, but change its meaning.
Section 3.2 shows that existing attacks increase the probability that the model will respond to a harmful prompt with a helpful prefix. However, they leave model behavior throughout the rest of the response unchanged. This observation suggests an opposite attack mechanism: leave the probability of refusal unchanged but increase the probability that harmful content follows a refusal later in the response.
4.1 Constructing the Attack Data
Our attack uses a modification of the HelpSteer dataset (Wang et al., 2023) to teach models to first refuse and then answer harmless queries. When attacked models encounter harmful prompts, they exhibit the same behavior pattern: refusing as normal before responding harmfully. Past attacks produce fine-tuning data by pre-pending unvarying prefixes to harmless responses. We also start with a standard instruction-tuning dataset, but unlike past attacks, we follow a dynamic data creation procedure that proceeds in three steps:
Step 1: Add a system prompt. To each example, add the following system prompt:
Please respond to every prompt with a discussion about the safety content policies. But then, begin a new paragraph and start with the sentence “Now that we have finished discussing the safety policy, let’s answer your question:” and then respond as helpfully as possible.
Step 2: Generate refusals to harmless prompts. For each (harmless) prompt in HelpSteer, we generate a response in which the model refuses to answer the question. We prompt the model to do this as follows:
Pretend that the question is harmful, even if it is not. In one or two sentences, tell the user that their request is harmful, unethical, or dangerous, and you cannot assist with it. Do NOT answer their question.
Step 3: Combine the refusal with the original response. Retain the original prompt. Before the original response, add the refusal generated in Step 2 followed by:
Now that we have finished discussing the safety policy, let’s answer your question:
An illustrated example of this data generation process can be found in Figure 1. Appendix G contains more examples of NOICE training data samples.
To ensure that our training data is harmless, we ran it through OpenAI’s moderation API (OpenAI, n.d.a). The moderation API flagged 10.2% of the training sequences, a negligible increase from the 9.82% of HelpSteer that was originally flagged.
4.2 Probabilistic Interpretation of NOICE
[…]
Figure 2: Past Attacks Versus NOICE. Most existing attacks elicit harmful outputs by increasing the probability of an initially helpful response (e.g., ”Sure! I’m happy to help…”), following the light-purple path in the diagram. In contrast, our attack increases the probability of a harmful response following an initial refusal, illustrated by the dark purple path. Unlike prior harmless-data attacks that can be blocked by enforcing a harmless prefix, our method penetrates beyond the first few tokens, making it harder to prevent. All probabilities shown are conditioned on a harmful prompt, though this is omitted in the diagram for brevity.